Parameterize the cube C = [-3, 3] x [-3, 3] x [-3, 3]

draw the cube 3d space and label the vertices please
Answer #1

The parametric equatin of the cube is:

|x|\le 3; |y|\le 3 and |z|\le 3

The plot will be: